Determine the derivative of "y = e" cos2"^x"


1
Expert's answer
2021-06-12T00:22:50-0400

Let us determine the derivative of "y = e^{\cos 2x}" using chain rule "(f(g(x)))'=f'(g(x))\cdot g'(x)."


"y'=e^{\cos 2x}(\cos 2x)'=e^{\cos 2x}(-\sin 2x)(2x)'=-2\sin 2xe^{\cos 2x}."